# Competitive Equilibrium for Electricity Markets with Spatially Flexible Loads

### v2

@@ Abstract
Electricity market, Datacenters, Electric Vehicles, Spatial Flexibility.
@@ I Introduction
With the electrification of transportation and the rapid expansion of large-scale, energy-intensive computing infrastructures for AI, power systems are becoming increasingly complex and more tightly integrated with other sectors. In 2023, electric vehicle (EV) electricity consumption in the U.S. surged to 7,596 GWh, nearly five times the level in 2018 [1]. Simultaneously, datacenters accounted for approximately 4% of U.S. electricity demand in 2022, a share projected to reach nearly 6% by 2026 [2]. However, the impact of these new loads extends far beyond this sheer increase in demand. Their nature differs fundamentally from conventional demand, which is fixed at a single location. In contrast, both EVs and geo-distributed datacenters can shift their energy consumption across different locations and this distinctive characteristic defines them as spatially flexible loads (FLs). Furthermore, the collective FLs may depend on external infrastructure networks, such as the transportation network for EVs and compute/communication infrastructure for datacenters, leading to the notion of FL systems that interconnect with the grid.
@@ I-B Related Literature
Studies on power-transportation coupling generally fall into two categories. The first assumes a central coordinator, which minimizes combined transportation and power system costs [3], guides EV routing to alleviate grid congestion [4], or reduces overall operational costs of EVs [5]. The second extends classical transportation models by considering a decentralized setting, where numerous EV owners individually select routes and charging locations to minimize travel costs, thereby influencing grid operation. While some studies omit charging costs in route decisions [6] or treat charging prices as exogenous retail rates [7], Alizadeh et al. [8] model EV charging within a competitive electricity market and capture the bidirectional coupling between LMPs and EV behavior, a setting conceptually closest to ours. This model is also applied in [9, 10] to the distribution power system studies. Building on these models, subsequent research designs charging pricing mechanisms to reduce grid peaks [11], maximize charging station profitability [12, 13], or align user behavior with social welfare [14, 15, 16].
Research on datacenter-power system coupling has largely been unidirectional. The first direction models the grid’s impact on datacenters, focusing on how LMPs influence datacenter workload allocation across geographical locations to minimize electricity costs [17, 18, 19] or carbon emissions [20, 21]. The second direction models the datacenter’s impact on the grid, particularly how spatial flexibility aids demand response [22, 23, 24], or affects market outcomes [25, 26]. While a few recent studies address bidirectional interactions, they generally operate outside the framework of competitive wholesale markets. These studies employ price signals other than LMPs to model the datacenter’s response, including Stackelberg game formulations [27], or auction-based market designs [28] and flexibility pricing schemes [29].
@@ II-A Models of Generators and Stationary Loads
We consider a power system in which each bus is equipped with a generator and a stationary load. The generation at bus $i$ is denoted by $g_{i}$, with feasible set $\mathcal{G}_{i}\subseteq\mathbb{R}$ being non-empty, compact and convex. Let $\mathbf{g}\in\mathbb{R}^{N}$ denote the vector of all generator outputs. Each generator optimizes its output in response to the LMP $\lambda_{i}$ at its bus. Assuming a stictly convex cost function $c_{i}(g_{i})$, the cost-minimization problem for the generator at bus $i$ is given by:
@@
Meanwhile, each bus also hosts a stationary load with consumption level $\ell_{i}$, which is assumed to be inelastic; let $\bm{\ell}\in\mathbb{R}^{N}$ denote the vector of all such loads.
@@ II-D Classical and Generalized Competitive Equilibrium
CE in electricity markets is well-studied and is known to exist under standard convexity assumptions. Moreover, it yields a socially optimal allocation and enjoys uniqueness under appropriate conditions [30, 31]. Building on these classical insights, we aim to investigate whether/when these desirable properties extend to the GCE framework, which is formalized and analyzed in the following sections.
@@ III-A Grid Coupled with Single FL System
Theorem 1 indicates that the closed-loop feedback between the grid and the FL system can be characterized by a convex optimization problem, and the equilibrium price $\bm{\lambda}^{\star}$ coincides with the Lagrange multiplier of the nodal power-balance constraint in that optimization. Under the scenario when $J(\mathbf{s})$ is available or can be learned from data, ($\mathrm{P}_{\mathrm{GCE}}$) allows us to abstract away individual FL decisions while preserving their aggregate response to LMPs. This formulation captures the collective behavior of a large population of FL through a low-dimensional, system-level representation.
@@ III-B Generalization to Multiple Independent FL Systems
Consider multiple FL systems connected to the same power system, each responding to common LMPs but without direct interaction beyond their shared grid connection, and each satisfying the framework in Section II. Assuming there are $W$ FL systems, we begin by redefining the variables: Let $\mathbf{S}\in\mathbb{R}^{N\times W}$ denote the matrix of electricity consumption, and its $w$-th column $\mathbf{S}_{w}$ represents the load vector of the $w$-th FL system, associated with its own feasible set $\mathcal{S}_{w}$ and value functions $J_{w}(\cdot)$ and $J_{w}^{\mathrm{SW}}(\cdot)$.
@@ IV-A Elements of the FL System for Electrified Transportation
We model the transportation network as a directed graph $\mathcal{G}^{\mathrm{tr}}=(\mathcal{V}^{\mathrm{tr}},\mathcal{E}^{\mathrm{tr}})$, where $\mathcal{V}^{\mathrm{tr}}$ denotes the set of locations (a subset of which host EV charging stations) and $\mathcal{E}^{\mathrm{tr}}$ denotes the set of directed roads connecting these locations. Let $N_{\mathrm{C}}$ denote the number of charging stations. Let there be $K$ origin-destination (OD) pairs, i.e., $\{(o_{1},d_{1}),\ldots,(o_{K},d_{K})\}$ with $(o_{k},d_{k})\in\mathcal{V}^{\mathrm{tr}}\times\mathcal{V}^{\mathrm{tr}}$. For each OD pair $k$, let $\mathcal{R}_{k}^{\mathrm{tr}}$ denote the finite set of feasible routes, each route being an ordered list of edges. Define the set of all routes as $\mathcal{R}^{\mathrm{tr}}=\bigcup_{k=1}^{K}\mathcal{R}_{k}^{\mathrm{tr}}$ and the number of routes as $N_{\mathrm{R}}=|\mathcal{R}^{\mathrm{tr}}|$. Each route is assumed to include exactly one charging station. [^1] If a route contains multiple charging stations, we conceptually split it into distinct routes, each associated with a single charging station along the path. Define the charger-route incidence matrix $\mathbf{A}^{\mathrm{CR}}\in\mathbb{R}^{N_{\mathrm{C}}\times N_{\mathrm{R}}}$, whose entry $(j,r)$ is 1 if the $r$-th route includes the $j$-th charging station, and 0 otherwise. Each column of $\mathbf{A}^{\mathrm{CR}}$ contains exactly one nonzero entry. Similarly, define the charger-bus incidence matrix $\mathbf{A}^{\mathrm{CB}}\in\mathbb{R}^{N_{\mathrm{C}}\times N}$, where entry $(j,i)$ is 1 if the $j$-th charging station is connected to bus $i$. Denote the total number of EVs in the transportation system by $N_{\mathrm{EV}}$. For simplicity, each EV is assumed to charge a fixed amount of energy $q>0$.
@@ IV-A2 FL decisions
Given the traffic flow $\mathbf{x}$}, the aggregate charging load at each power system bus satisfies (2), where matrix $\mathbf{A}^{\mathrm{FL}}$ is given by $\mathbf{A}^{\mathrm{FL}}={q}\cdot(\mathbf{A}^{\mathrm{CB}})^{\top}\mathbf{A}^{\mathrm{CR}}$.
@@ IV-A4 FL best response model
We are interested in characterizing how the LMPs affect EVs’ route and charging location choices. Given the LMPs, each EV driver selects a route and charging station aiming to minimize their individual cost. The collective outcome of these selfish individual decisions is characterized by the Nash equilibrium of the nonatomic game, commonly referred to as the user equilibrium (UE) in the transportation literature [32]. Specifically, given the LMP $\bm{\lambda}$, the costs associated with a route $R_{r}$ is decomposed into travel and charging components:
@@ V-A Elements of the FL System
Suppose there are $N_{\mathrm{D}}$ datacenters in the system, and each datacenter is connected to a power grid bus. We define the datacenter-bus incidence matrix as $\mathbf{A}^{\mathrm{DB}}\in\mathbb{R}^{N_{\mathrm{D}}\times N}$, where entry $(j,i)$ is 1 if the $j$-th charging station is connected to bus $i$ in the power grid. Each unit of computing workload consumes a fixed amount of electricity $q>0$.
@@ V-A4 FL best response model
We assume that each AI company directly pays for the electricity consumed by its workloads in datacenters, an arrangement consistent with self-operated or colocation datacenters, and serving as a stylized abstraction of electricity bill pass-through in cloud computing environments. The net payoff of company $k$ is then the utility from computing workloads minus the cost of electricity. In response to the LMPs $\bm{\lambda}$, each company solves:
@@ V-B Structural Results for Coupled FL and Power System
The term $\tfrac{\partial}{\partial\mathbf{x}^{(\tilde{k})}}\!\left(\nabla_{\mathbf{x}^{(k)}}u_{k}\right)$ represent the Jacobian matrix of the gradient of player $k$’s utility with respect to another player’s strategy, representing the cross-partial derivatives $\tfrac{\partial^{2}u_{k}}{\partial\mathbf{x}^{(\tilde{k})}\partial\mathbf{x}^{(k)}}$. The symmetry cross-partial-derivative condition (22) ensures that the game is a potential game [33], permitting the existence of a potential function $\Phi(\mathbf{x})$. This condition frequently holds in datacenter environments, as companies sharing computational resources commonly experience reciprocal performance effects, such as shared queueing delays, leading to symmetric interactions in their utility functions. Concave utility functions $u_{k}$ are also typical in practice, as the marginal value of computing resources generally diminishes with increasing workload. Given the conditions in Lemma 4, Assumption 2 holds for datacenter networks. Given the definition of $\Phi(\mathbf{x})$, the corresponding value function $J(\mathbf{s})$ follows from (4), and the optimization-based characterization for the GCE in Theorem 1 then applies to the setting here. If the potential-game condition (22) is not satisfied, a merit-function VI approach is used; see Appendix D.
@@ V-C Illustrative Example
As illustrated in Fig. 3, the system consists of two AI companies that allocate their computing workloads across two data centers to optimize their utility, i.e., $K=2$. In this setup, the two data centers are connected to power system buses 1 and 2, respectively. We do not enforce capacity limits on the datacenter loads, allowing them to be flexible in their load-shifting strategies. Assume there is no stationary demand at the buses, and the only load is from the datacenters. Consequently, the total demand at bus 1 is $s_{1}=q\big(x_{1}^{(1)}+x_{1}^{(2)}\big)$, and the total demand at bus 2 is $s_{2}=q\big(x_{2}^{(1)}+x_{2}^{(2)}\big)$. Additionally, two generators at bus 1 and bus 2 provide power outputs $g_{1}$ and $g_{2}$, respectively, with quadratic cost function $\gamma_{i}g_{i}^{2}$, $i=1,2$.
In this example, we assume the utility function from computing workload for company $k$, $u_{k}\left(\mathbf{x}^{(k)};\mathbf{x}^{(-k)}\right)$, is quadratic:
@@
where $\bm{\eta}_{k}\in\mathbb{R}^{N_{\mathrm{D}}}$ denotes the baseline utility, reflecting factors such as locational preference and latency penalties associated with geographical distance between datacenters and companies. The symmetric matrix $\mathbf{M}_{k,k}\in\mathbb{R}^{N_{\mathrm{D}}\times N_{\mathrm{D}}}$ captures diminishing returns to scale in resource utilization across datacenters, while $\mathbf{M}_{k,-k}\in\mathbb{R}^{N_{\mathrm{D}}\times(K-1)N_{\mathrm{D}}}$ represents the influence of other companies’ workloads on company $k$’s utility, including queueing delays at each datacenter [34].
@@
Under the symmetric condition (22), $\mathbf{M}^{\mathrm{NE}}$ is required to be symmetric, i.e., $\mathbf{M}^{\mathrm{NE}}=(\mathbf{M}^{\mathrm{NE}})^{\top}$, and the corresponding FL preference function $\Phi(\mathbf{x})$ is given by

### v3

@@ Abstract
Datacenters, Electricity market, Electric vehicles, Spatial flexibility.
@@ I Introduction
With the electrification of transportation and the rapid expansion of large-scale, energy-intensive computing infrastructures for AI, power systems are becoming increasingly complex and more tightly integrated with other sectors. In 2023, electric vehicle (EV) electricity consumption in the U.S. surged to 7,596 GWh, nearly five times the level in 2018 [ev-increase]. Simultaneously, datacenters accounted for approximately 4% of U.S. electricity demand in 2022, a share projected to reach nearly 6% by 2026 [dc-increase]. However, the impact of these new loads extends far beyond this sheer increase in demand. Their nature differs fundamentally from conventional demand, which is fixed at a single location. In contrast, both EVs and geo-distributed datacenters can shift their energy consumption across different locations and this distinctive characteristic defines them as spatially flexible loads (FLs). Furthermore, the collective FLs may depend on external infrastructure networks, such as the transportation network for EVs and compute/communication infrastructure for datacenters, leading to the notion of FL systems that interconnect with the grid.
@@ I-B Related Literature
Studies on power-transportation coupling generally fall into two categories. The first assumes a central coordinator, which minimizes combined transportation and power system costs [zhang2020power], guides EV routing to alleviate grid congestion [guo2014rapid], or reduces overall operational costs of EVs [8737720]. The second extends classical transportation models by considering a decentralized setting, where numerous EV owners individually select routes and charging locations to minimize travel costs, thereby influencing grid operation. While some studies omit charging costs in route decisions [8445618] or treat charging prices as exogenous retail rates [7572952], Alizadeh et al. [alizadeh2016optimal] model EV charging within a competitive electricity market and capture the bidirectional coupling between LMPs and EV behavior, a setting conceptually closest to ours. This model is also applied in [7967870, 9891825] to the distribution power system studies. Building on these models, subsequent research designs charging pricing mechanisms to reduce grid peaks [tan2015real], maximize charging station profitability [9330803, 8304623], or align user behavior with social welfare [10644866, sheng2021coordinated, 9748107].
Research on datacenter-power system coupling has largely been unidirectional. The first direction models the grid’s impact on datacenters, focusing on how LMPs influence datacenter workload allocation across geographical locations to minimize electricity costs [li2024towards, 6322266, 9963634] or carbon emissions [hall2024carbon, 9770383]. The second direction models the datacenter’s impact on the grid, particularly how spatial flexibility aids demand response [wierman2014opportunities, fridgen2017shifting, 7517380], or affects market outcomes [ZHANG2020106723, zhang2022remunerating]. While a few recent studies address bidirectional interactions, they generally operate outside the framework of competitive wholesale markets. These studies employ price signals other than LMPs to model the datacenter’s response, including Stackelberg game formulations [10636017], or auction-based market designs [8540414] and flexibility pricing schemes [werner2021pricing].
@@ II-A Models of Generators and Stationary Loads
We consider a power system in which each bus is equipped with a generator and a stationary load. The generation at bus $i$ is denoted by $g_{i}$, with feasible set $\mathcal{G}_{i}\subseteq\mathbb{R}$ being non-empty, compact and convex. Let $\mathbf{g}\in\mathbb{R}^{N}$ denote the vector of all generator outputs. Each generator optimizes its output in response to the LMP $\lambda_{i}$ at its bus. With a strictly convex cost function $c_{i}(g_{i})$, the cost-minimization problem for the generator at bus $i$ is given by:
@@
Meanwhile, each bus also hosts a stationary load with consumption level $\ell_{i}$, which is modeled to be inelastic; let $\bm{\ell}\in\mathbb{R}^{N}$ denote the vector of all such loads.
@@ II-B Model of Spatially Flexible Loads
Remark 1 (Interpreting $\mathbf{x}$ and Justifying Assumption 1).
The vector $\mathbf{x}\in\mathcal{X}$ collects the internal decision variables of the FL system, i.e., the representation of how flexible demand is allocated across its own network (e.g., route/charging choices in a transportation network, or workload allocation across datacenter sites). Each component $x_{k}$ is a continuous variable that measures the amount or intensity of decision option $k$; it may be inherently continuous or arise as a continuous abstraction or relaxation of an underlying physical quantity such as counts or flows. The matrix $\mathbf{A}^{\mathrm{FL}}$ encodes how these internal decisions map into nodal net consumption as observed by the power system. The linear form (2) reflects a general aggregation principle: each decision option $x_{k}$ is geographically associated with a specific bus at which it draws power, while multiple options may aggregate at the same bus. The coefficient $\mathbf{A}^{\mathrm{FL}}]_{ik}$ captures/approximates the fixed or average power requirement [Desislavov_2023] per unit of $x_{k}$ at bus $i$. Heterogeneous power requirements can be accommodated either by allowing these coefficients to vary across options or by refining the decision set $\mathbf{x}$ into more granular categories. In Sections IV and V, we specify $\mathbf{x}$ and $\mathbf{A}^{\mathrm{FL}}$ explicitly for EVs and datacenters.
@@
Remark 2 (Interpreting $\Phi(\mathbf{x})$ and Justifying Assumption 2).
The optimization form in (3) is a behavioral representation and does not necessarily require the existence of a central controller. Its general form can encapsulate two distinct classes of settings: (i) Centralized Coordination, where a coordinator directly minimizes a system-level FL objective, and (ii) Decentralized Equilibrium, where many self-interested agents interact (through mechanisms like competition for shared resources) and the induced Nash equilibrium coincides with the minimizer of a potential function $\Phi(\mathbf{x})+\bm{\lambda}^{\top}\mathbf{A}^{\mathrm{FL}}\mathbf{x}$ (non-potential cases are addressed in Lemma 4 and Appendix A). In both settings, $\Phi(\mathbf{x})$ encodes preferences that are internal to the FL system and independent of electricity prices. The term $\bm{\lambda}^{\top}\mathbf{A}^{\mathrm{FL}}\mathbf{x}$ represents the electricity payment by the FL system to the grid. Sections IV and V detail $\Phi(\mathbf{x})$ for specific sectors.
@@ II-C Economic Dispatch
where $\mathbf{p}\in\mathbb{R}^{N}$ represents the vector of power injections at each bus. The feasible set $\mathcal{P}\subset\mathbb{R}^{N}$ is defined under the standard DC power flow model as a convex and compact polyhedral set: $\mathcal{P}:=\left\{\mathbf{p}\in\mathbb{R}^{N}\;\middle|\;\mathbf{1}^{\top}\mathbf{p}=0,\;-{\mathbf{f}}^{\max}\leq\mathbf{H}\mathbf{p}\leq{\mathbf{f}^{\max}}\right\},$ where the first condition ensures system-wide power balance, and the second enforces line flow limits using the shift-factor matrix $\mathbf{H}\in\mathbb{R}^{M\times N}$, with ${\mathbf{f}^{\max}}\in\mathbb{R}^{M}$ denoting the vector of line capacity bounds. The dual variable $\bm{\lambda}$ on the nodal power balance constraint corresponds to the LMPs.
Remark 3 (Demand Modeling in Economic Dispatch).
@@ II-D Classical and Generalized Competitive Equilibrium
CE in electricity markets is well-studied and is known to exist under standard convexity assumptions. Moreover, it yields a socially optimal allocation and enjoys uniqueness under appropriate conditions [hsu1997introduction, 4111733]. Building on these classical insights, we aim to investigate whether/when these desirable properties extend to the GCE framework, which is formalized and analyzed in the following sections.
@@ III-A Grid Coupled with Single FL System
Theorem 1 indicates that the closed-loop feedback between the grid and the FL system can be characterized by a convex optimization problem, and the equilibrium price $\bm{\lambda}^{\star}$ coincides with the Lagrange multiplier of the nodal power balance constraint in that optimization. Under the scenario when $J(\mathbf{s})$ is available or can be learned from data, ($\mathrm{P}_{\mathrm{GCE}}$) allows us to abstract away individual FL decisions while preserving their aggregate response to LMPs. This formulation captures the collective behavior of a large population of FL through a low-dimensional, system-level representation.
@@ III-B Generalization to Multiple Independent FL Systems
Consider multiple FL systems connected to the same power system, each responding to common LMPs but without direct interaction beyond their shared grid connection, and each satisfying the framework in Section II. Consider a setting with $W$ FL systems. We begin by redefining the variables: Let $\mathbf{S}\in\mathbb{R}^{N\times W}$ denote the matrix of electricity consumption, and its $w$-th column $\mathbf{S}_{w}$ represents the load vector of the $w$-th FL system, associated with its own feasible set $\mathcal{S}_{w}$ and value functions $J_{w}(\cdot)$ and $J_{w}^{\mathrm{SW}}(\cdot)$.
@@ IV-A Elements of the FL System for Electrified Transportation
We model the transportation network as a directed graph $\mathcal{G}^{\mathrm{tr}}=(\mathcal{V}^{\mathrm{tr}},\mathcal{E}^{\mathrm{tr}})$, where $\mathcal{V}^{\mathrm{tr}}$ denotes the set of locations (a subset of which host EV charging stations) and $\mathcal{E}^{\mathrm{tr}}$ denotes the set of directed roads connecting these locations. Let $N_{\mathrm{C}}$ denote the number of charging stations. Let there be $K$ origin-destination (OD) pairs, i.e., $\{(o_{1},d_{1}),\ldots,(o_{K},d_{K})\}$ with $(o_{k},d_{k})\in\mathcal{V}^{\mathrm{tr}}\times\mathcal{V}^{\mathrm{tr}}$. For each OD pair $k$, let $\mathcal{R}_{k}^{\mathrm{tr}}$ denote the finite set of feasible routes, each route being an ordered list of edges. Define the set of all routes as $\mathcal{R}^{\mathrm{tr}}=\bigcup_{k=1}^{K}\mathcal{R}_{k}^{\mathrm{tr}}$ and the number of routes as $N_{\mathrm{R}}=|\mathcal{R}^{\mathrm{tr}}|$. Each route is assumed to include exactly one charging station. [^1] If a route contains multiple charging stations, we conceptually split it into distinct routes, each associated with a single charging station along the path. Define the charger-route incidence matrix $\mathbf{A}^{\mathrm{CR}}\in\mathbb{R}^{N_{\mathrm{C}}\times N_{\mathrm{R}}}$, whose entry $(j,r)$ is 1 if the $r$-th route includes the $j$-th charging station, and 0 otherwise. Each column of $\mathbf{A}^{\mathrm{CR}}$ contains exactly one nonzero entry. Similarly, define the charger-bus incidence matrix $\mathbf{A}^{\mathrm{CB}}\in\mathbb{R}^{N_{\mathrm{C}}\times N}$, where entry $(j,i)$ is 1 if the $j$-th charging station is connected to bus $i$. Denote the total number of EVs in the transportation system by $N_{\mathrm{EV}}$.
@@ IV-A2 FL decisions
For simplicity, each EV is assumed to charge a fixed amount of energy $q>0$. Given the traffic flow $\mathbf{x}$}, the aggregate charging load at each power system bus satisfies (2), where matrix $\mathbf{A}^{\mathrm{FL}}$ is given by $\mathbf{A}^{\mathrm{FL}}={q}\cdot(\mathbf{A}^{\mathrm{CB}})^{\top}\mathbf{A}^{\mathrm{CR}}$.
Remark 4 (Heterogeneous Charging Energy Requirements).
If EVs have heterogeneous charging energy requirements, partition them into $\Upsilon$ categories with per-EV energy $q_{\upsilon}$. Let $\mathbf{x}^{(\upsilon)}\in\mathbb{R}_{+}^{N_{\mathrm{R}}}$ be the route traffic flow vector for category $\upsilon$ and stack $\mathbf{x}:=[(\mathbf{x}^{(1)})^{\top},\ldots,(\mathbf{x}^{(\Upsilon)})^{\top}]^{\top}$. Then Assumption 1 still holds with matrix $\mathbf{A}^{\mathrm{FL}}$ given by $\mathbf{A}^{\mathrm{FL}}=[q_{1}(\mathbf{A}^{\mathrm{CB}})^{\top}\mathbf{A}^{\mathrm{CR}}\ \cdots\ q_{\Upsilon}(\mathbf{A}^{\mathrm{CB}})^{\top}\mathbf{A}^{\mathrm{CR}}].$
@@ IV-A4 FL best response model
We are interested in characterizing how the LMPs affect EVs’ route and charging location choices. Given the LMPs, each EV driver selects a route and charging station aiming to minimize their individual cost. The collective outcome of these selfish individual decisions is characterized by the Nash equilibrium of the nonatomic game, commonly referred to as the user equilibrium (UE) in the transportation literature [sheffi1985urban]. Specifically, given the LMP $\bm{\lambda}$, the costs associated with a route $R_{r}$ is decomposed into travel and charging components:
@@ V-A Elements of the FL System
Suppose there are $N_{\mathrm{D}}$ datacenters in the system, and each datacenter is connected to a power grid bus. We define the datacenter-bus incidence matrix as $\mathbf{A}^{\mathrm{DB}}\in\mathbb{R}^{N_{\mathrm{D}}\times N}$, where entry $(j,i)$ is 1 if the $j$-th datacenter is connected to bus $i$ in the power grid. Each unit of computing workload consumes a fixed amount of electricity $q>0$ [Desislavov_2023].
@@ V-A4 FL best response model
In our model, each AI company directly pays for the electricity consumed by its workloads in datacenters, an arrangement consistent with self-operated or colocation datacenters, and serving as a stylized abstraction of electricity bill pass-through in cloud computing environments. The net payoff of company $k$ is then the utility from computing workloads minus the cost of electricity. In response to the LMPs $\bm{\lambda}$, each company solves:
@@ V-B Structural Results for Coupled FL and Power System
The term $\tfrac{\partial}{\partial\mathbf{x}^{(\tilde{k})}}\!\left(\nabla_{\mathbf{x}^{(k)}}u_{k}\right)$ represent the Jacobian matrix of the gradient of player $k$’s utility with respect to another player’s strategy, representing the cross-partial derivatives $\tfrac{\partial^{2}u_{k}}{\partial\mathbf{x}^{(\tilde{k})}\partial\mathbf{x}^{(k)}}$. The symmetry cross-partial-derivative condition (22) ensures that the game is a potential game [monderer1996potential], permitting the existence of a potential function $\Phi(\mathbf{x})$. This condition frequently holds in datacenter environments, as companies sharing computational resources commonly experience reciprocal performance effects, such as shared queueing delays, leading to symmetric interactions in their utility functions. Concave utility functions $u_{k}$ are also typical in practice, as the marginal value of computing resources generally diminishes with increasing workload. Given the conditions in Lemma 4, Assumption 2 holds for datacenter networks. Given the definition of $\Phi(\mathbf{x})$, the corresponding value function $J(\mathbf{s})$ follows from (4), and the optimization-based characterization for the GCE in Theorem 1 then applies to the setting here. If the potential-game condition (22) is not satisfied, a merit-function VI approach is used; see Appendix A.
@@ V-C Illustrative Example
As illustrated in Fig. 3, the system consists of two AI companies that allocate their computing workloads across two data centers to optimize their utility, i.e., $K=2$. In this setup, the two data centers are connected to power system buses 1 and 2, respectively. We do not enforce capacity limits on the datacenter loads, allowing them to be flexible in their load-shifting strategies. There is no stationary demand at the buses; all load comes from the datacenters. Consequently, the total demand at bus 1 is $s_{1}=q\big(x_{1}^{(1)}+x_{1}^{(2)}\big)$, and the total demand at bus 2 is $s_{2}=q\big(x_{2}^{(1)}+x_{2}^{(2)}\big)$. Additionally, two generators at bus 1 and bus 2 provide power outputs $g_{1}$ and $g_{2}$, respectively, with quadratic cost function $\gamma_{i}g_{i}^{2}$, $i=1,2$.
In this example, we set the utility function from computing workload for company $k$, $u_{k}\left(\mathbf{x}^{(k)};\mathbf{x}^{(-k)}\right)$, is quadratic:
@@
where $\bm{\eta}_{k}\in\mathbb{R}^{N_{\mathrm{D}}}$ denotes the baseline utility, reflecting factors such as locational preference and latency penalties associated with geographical distance between datacenters and companies. The symmetric matrix $\mathbf{M}_{k,k}\in\mathbb{R}^{N_{\mathrm{D}}\times N_{\mathrm{D}}}$ captures diminishing returns to scale in resource utilization across datacenters, while $\mathbf{M}_{k,-k}\in\mathbb{R}^{N_{\mathrm{D}}\times(K-1)N_{\mathrm{D}}}$ represents the influence of other companies’ workloads on company $k$’s utility, including queueing delays at each datacenter [tsiligkaridis2025distributed].
@@
Under the symmetric condition (22), $\mathbf{M}^{\mathrm{NE}}$ is required to be symmetric, i.e., $\mathbf{M}^{\mathrm{NE}}=(\mathbf{M}^{\mathrm{NE}})^{\top}$. According to Lemma 4, the corresponding FL preference function $\Phi(\mathbf{x})$ is given by